The lifetime of a machine is continuous on the interval (0, 40) with probability density function f, where f(t) is proportional to (t + 10)−2, and t is the lifetime in years. Calculate the probability that the lifetime of the machine part is less than 10 years. Hint: Show that f(t) is legitimate and find the proportionality constant.
1
Expert's answer
2020-05-15T17:55:22-0400

Let "c=const," then


"f(t) = \begin{cases}
 c(t+10)^{-2} & 0\leq t\leq 40 \\
 0 &otherwise
\end{cases}"

"1=\displaystyle\int_{-\infin}^{\infin}f(t)dt=\displaystyle\int_{0}^{40}{c\over (t+10)^2}dt="


"=-c\bigg[{1\over t+10}\bigg]\begin{matrix}
 40 \\
 0
\end{matrix}=-c\bigg({1\over 40+10}-{1\over 0+10}\bigg)=0.08c"

"c=12.5"

"f(t) = \begin{cases}
 12.5(t+10)^{-2} & 0\leq t\leq 40 \\
 0 &otherwise
\end{cases}"

Calculate the probability that the lifetime of the machine part is less than 10 years. 


"P(T<10)=\displaystyle\int_{-\infin}^{10}f(t)dt=\displaystyle\int_{0}^{10}{12.5\over (t+10)^2}dt="

"=-12.5\bigg[{1\over t+10}\bigg]\begin{matrix}
 10 \\
 0
\end{matrix}=-12.5\bigg({1\over 10+10}-{1\over 0+10}\bigg)=0.625"